# Semantic Self-Distillation for Language Model Uncertainty

### v1

@@ 3.1 Models and Dataset Generation
We evaluate the SSD distribution in the context of short-form question answering, using the TriviaQA dataset (Joshi et al., 2017). We construct a training set of $N_{train}=4000$ prompts and report results on a held-out test set of $N_{test}=1000$ prompts.
We evaluate across five model families: Mistral (Ministral-8B-Instruct), Llama-3 (3.1-8B-Instruct, 3.2-3B-Instruct), Qwen (Qwen3-4B-Instruct-2507, Qwen3-8B), Gemma (Gemma-3-4B) and SmolLM (SmolLM3-3B). For all experiments, we use Qwen3-8B as the oracle ‘judge’ model to determine the correctness of generated answers.
@@
For each prompt and model, we sample $S=32$ samples at temperature $T=1$. We additionally sample a single ‘default’ answer at $T=0.1$ which we input to the judge model to generate the hallucination label.
@@
For the answer embedding, we use EmbeddingGemma (Vera et al., 2025), with the embedding truncated to a dimension of $d_{z}=128$ . For hallucination detection, we further reduce dimensionality using PCA to $d_{z}^{\prime}=16$, following ablation studies detailed in Appendix Section D. For likelihood-based OOD verification, we set $d_{z}^{\prime}=32$.
3.2 Layer Selection Protocol
To estimate a suitable layer $\ell^{*}$ for the student input, we use the results of our hallucination detection baseline experiments, described in Section 3.3. Following Kossen et al. (2024), we first train simple linear probes on all layers to predict the binarized Semantic Entropy. We select the layer index $\ell^{*}$ that maximizes validation accuracy for this proxy task. We assume that the layer most predictive of the scalar entropy value also contains the richest information for constructing the full distributional mixture. We find this typically corresponds to deeper layers, shown in Table 5 in the Appendix.
3.3 Hallucination Detection via Predicted Dispersion
@@
We treat the student’s dispersion score as a pre-generation risk score: higher values indicate greater semantic spread and higher hallucination risk. We report AUROC and AUPRC against correctness labels determined using an LLM-as-judge framework.
@@
We first compare the student’s dispersion score to the standard deviation of the teacher’s $S=32$ semantic embeddings, which we term teacher dispersion (TD). Table 1 reports results across models. Despite requiring no inference-time sampling, SSD outperforms TD on four of the seven models investigated in AUROC and on all seven in AUPRC. We attribute this in part to a smoothing effect: TD is a finite-sample statistic and assigns zero uncertainty to prompts where the model consistently produces the same answer, even when that answer is incorrect. By contrast, SSD learns a continuous mapping from prompt representations to semantic uncertainty, enabling differentiated risk estimates in such degenerate cases, which likely contributes to its stronger AUPRC.
Model AUROC AUPRC
TD SSD TD SSD
Qwen3-8B 0.684 $\pm$ 0.017 0.677 $\pm$ 0.018 0.525 $\pm$ 0.027 0.570 $\pm$ 0.028
Qwen3-4B 0.732 $\pm$ 0.016 0.699 $\pm$ 0.016 0.596 $\pm$ 0.025 0.618 $\pm$ 0.025
Llama 3.1 8B 0.695 $\pm$ 0.017 0.686 $\pm$ 0.020 0.443 $\pm$ 0.028 0.519 $\pm$ 0.030
Llama 3.2 3B 0.665 $\pm$ 0.017 0.738 $\pm$ 0.017 0.558 $\pm$ 0.027 0.657 $\pm$ 0.027
Ministral 8B 0.743 $\pm$ 0.016 0.748 $\pm$ 0.016 0.579 $\pm$ 0.029 0.599 $\pm$ 0.029
SmolLM3 3B 0.698 $\pm$ 0.016 0.711 $\pm$ 0.016 0.569 $\pm$ 0.025 0.623 $\pm$ 0.026
Gemma 3 4B 0.676 $\pm$ 0.017 0.701 $\pm$ 0.017 0.632 $\pm$ 0.024 0.662 $\pm$ 0.025
Table 1: Distilling sampled dispersion into a single pass. TD computes dispersion from $S=32$ sampled semantic embeddings at inference time, while SSD predicts a prompt-conditioned semantic density and scores dispersion analytically. We report hallucination detection AUROC/AUPRC as mean $\pm$ standard deviation over 1,000 bootstrap resamples of the test set; bold indicates the better of TD vs. SSD per metric.
@@
We further compare SSD against three widely used hallucination detection methods:
  1. 1.
    Probability of Correctness Probe (PCP): A logistic regression classifier trained on the prompt representation to predict binary correctness. Originally proposed as ‘P(I know)’ by Kadavath et al. (2022), this method requires externally provided correctness labels.
  2. 2.
    Semantic Entropy (SE): A state-of-the-art sampling-based approach that measures semantic dispersion by clustering $S$ stochastic completions using a natural language inference (NLI) model to assess mutual entailment (Farquhar et al., 2024).
  3. 3.
    Semantic Entropy Probe (SEP): A lightweight probe with the same architecture as PCP, trained to regress a binarized SE target, inheriting SE’s sampling and NLI-based target construction cost (Kossen et al., 2024).
@@
To reflect agentic deployment constraints, we report hallucination detection performance alongside method characteristics: inference latency, supervision target construction cost, and whether a method supports post-generation (posterior) signals in addition to pre-generation (prior) uncertainty. Table 2 summarizes these trade-offs and reports macro-average AUROC across model families; full per-model AUROC/AUPRC tables are provided in the Appendix (tables 8 and 9).
Overall, SE achieves the strongest AUROC, consistent with prior work showing that explicit sampling of multiple reasoning paths yields high-quality uncertainty signals (Kadavath et al., 2022). Among single-pass methods, PCP and SEP outperform SSD as pre-generation hallucination predictors, confirming that the binary classification problem is simpler than distilling the geometry of a full semantic distribution. However, SSD occupies a distinct and practically useful regime: it matches probe-level inference latency, avoids entailment-based target construction, and uniquely provides posterior signals as well as prior risk scores. The post-generation signals enable verification and filtering mechanisms that are not available to scalar probes.
Method AUROC $\uparrow$ Latency $\downarrow$ Target cost $\downarrow$ Posterior
SE 0.804 $\pm$ 0.036 $S$ samples + $S^{2}$ NLI $\times$
SEP 0.759 $\pm$ 0.026 Single-pass $S$ samples + $S^{2}$ NLI $\times$
PCP 0.771 $\pm$ 0.021 Single-pass $C_{\text{CL}}$ $\times$
TD 0.699 $\pm$ 0.027 $S$ samples $\times$
SSD 0.708 $\pm$ 0.024 Single-pass $S$ samples + $S\,\Phi$ $\ell(y),\ \mathbb{E}[\mathbf{z}]$
Table 2: Hallucination detection trade-offs. AUROC is reported as a macro-average (mean $\pm$ std) across models. Latency summarizes inference-time requirements (single-pass vs $S$ sampled generations, with entailment-based SE additionally requiring $S^{2}$ natural language inference (NLI) comparisons). Target cost denotes the computation required to construct supervision targets during training; $C_{CL}$ is the cost of acquiring correctness labels. Only SSD explicitly provides post-generation posterior signals in addition to pre-generation uncertainty.
3.4 Distillation Fidelity vs. Detection Performance
Figure 2 plots hallucination detection AUROC against distillation fidelity across base models and student hyperparameter settings. We observe a strong positive relationship, with a cross-model Spearman correlation of $\rho_{\text{meta}}=0.65$, indicating that detection performance is limited by the student’s ability to learn the prompt-to-dispersion mapping. The failure cases, notably Qwen3-4B, exhibit near-zero $\rho_{\text{fidelity}}$, suggesting that the relevant uncertainty signal is not accessible from the chosen prompt representation.
Figure 2: Distillation fidelity drives detection performance. The X-axis shows distillation fidelity ($\rho_{\text{fidelity}}$): the Spearman correlation between the student’s predicted entropy and the teacher dispersion across test prompts. Multiple points per model correspond to different student hyperparameter settings. $\rho_{\text{meta}}$ is the Spearman correlation between the X and Y variables; it shows that models where the student fails to learn the distribution (low $\rho_{\text{fidelity}}$) result in poor detection.
3.5 Context Verification via Out-of-Domain Detection
@@
We evaluate the SSD density as a post-generation verification signal by measuring its ability to detect mismatched prompt-answer pairs. We frame this as an out-of-domain detection task: for each test prompt $x_{i}$, we pair its representation $\mathbf{h}_{i}$ with (i) the model’s default answer $y_{i}$ and (ii) a random answer $y_{j}$ drawn from a different prompt. Each candidate answer is embedded as $\mathbf{z}=\Phi(y)$ and scored using the student’s posterior log-likelihood $\log q_{\phi}(\mathbf{z}\mid\mathbf{h})$. We report AUROC for distinguishing matched from mismatched pairs.
@@
As shown in Table 3, the posterior likelihood is a strong context-verification signal, achieving AUROCs above 0.95 for most base models. This suggests that SSD can serve as a low-latency guardrail in agentic workflows, for example detecting cache poisoning (Chen et al., 2024) or filtering retrieved content in retrieval-augmented generation (Lewis et al., 2020).
The likelihood provides a modest signal for post-generation hallucination detection, demonstrated by AUROCs above 0.5. This result could be interpreted through the lens of Semantic Density (Qiu and Miikkulainen, 2024), which posits that reliable samples tend to lie in denser regions of a semantic predictive space. Rather than approximating this distribution via samples and associated sequence likelihoods, our approach explicitly models it.
Table 3: Utility of SSD posterior likelihood. Out-of-domain (OOD) detection reports AUROC for distinguishing an answer belonging to the prompt from a random answer. Hallucination detection compares AUROC for pre-generation entropy vs. post-generation likelihood. While likelihood is a strong OOD detector, entropy remains the superior signal for hallucination classification. Values report mean AUROC with bootstrap standard deviation over 1,000 resamples.
Model OOD Hallucination
Likelihood Entropy Likelihood
Qwen3 8B 0.983 $\pm$ 0.002 0.697 $\pm$ 0.017 0.571 $\pm$ 0.019
Qwen3 4B 0.984 $\pm$ 0.002 0.703 $\pm$ 0.016 0.555 $\pm$ 0.018
Llama 3.1 8B 0.989 $\pm$ 0.002 0.702 $\pm$ 0.019 0.578 $\pm$ 0.020
Llama 3.2 3B 0.980 $\pm$ 0.003 0.728 $\pm$ 0.016 0.612 $\pm$ 0.018
Ministral 8B 0.944 $\pm$ 0.005 0.758 $\pm$ 0.016 0.616 $\pm$ 0.018
SmolLM3 3B 0.988 $\pm$ 0.002 0.676 $\pm$ 0.018 0.563 $\pm$ 0.018
Gemma 3 4B 0.895 $\pm$ 0.007 0.679 $\pm$ 0.017 0.577 $\pm$ 0.018
3.6 Semantic Consensus from the Predicted Mixture
We evaluate the generative utility of the student by measuring its ability to approximate the semantic consensus of the model. We treat the empirical mean of $S=32$ stochastic samples as a reference semantic centroid, then measure the mean squared distance (MSD) in Euclidean space between this centroid and both (i) the model’s low-temperature output and (ii) the analytical mean of the student mixture, $\bm{\mu}_{ssd}=\sum_{k}\pi_{k}\bm{\mu}_{k}$. Lower MSD indicates better recovery of the model’s aggregate semantic belief. MSD is computed as $\mathbb{E}_{i}\left[\|\hat{\mathbf{z}}_{i}-\bm{\mu}^{\text{true}}_{i}\|_{2}^{2}\right]$ over examples $i$ in the subset.
We report results for the best-performing student configuration (Ministral 8B), reflecting the strong architectural dependence of distillation fidelity. As shown in Table 4, the SSD mixture mean provides a closer approximation to the semantic centroid than a single default decoding pass. This effect is most pronounced for incorrect answers, where SSD improves centroid recovery in over 76% of test cases. This aligns with prior work showing that hallucinated outputs tend to deviate from semantic consensus (Lin et al., 2024; Phillips et al., 2025). Together with the OOD results in Table 3, this result suggests that SSD’s pre-generation uncertainty estimates can be complemented by post-generation distance- and likelihood-based checks derived from the same predicted density.
Table 4: Semantic consensus approximation with Ministral 8B. We report the mean squared distance (MSD) of the default answer embedding and the SSD predicted mean to the true semantic centroid (average of $S=32$ samples). Values denote mean MSD, with the bootstrap standard deviation expressed as a percentage of the mean shown as a subscript.
All 0.0153.2 0.0081.9 +45.5 58.3
Correct 0.0134.4 0.0092.3 +32.3 48.8
Incorrect 0.0184.2 0.0073.4 +63.0 76.5
4 Future Work

### v2

@@ 4 Experiments
We evaluate SSD across multiple LLM families in order to answer two core questions. First, does SSD faithfully distil sampling-based semantic dispersion into a single-pass model? Second, what additional capabilities arise from modelling a full sequence-level density rather than a scalar uncertainty proxy? To address the first question, we study prompt-level hallucination prediction, treating the student’s entropy as a pre-generation risk score and comparing it against sampling-based and probing baselines. To address the second, we evaluate answer-level likelihood scores, demonstrating how the predicted density supports post-generation verification, including out-of-domain answer detection and multiple-choice answer selection. Figure 2 illustrates examples of each operation type.
@@ 4.1 Models and Dataset Generation
We evaluate SSD in the context of short-form question answering, using the TriviaQA [joshi-etal-2017-triviaqa] and MMLU [hendrycks2021measuring] datasets. We construct training sets of $N_{train}=4{,}000$ prompts and report results on held-out test sets of $N_{test}=1{,}000$ prompts.
MMLU is formatted as a multiple-choice benchmark, as opposed to the open-ended TriviaQA. To properly assess semantic dispersion and enable comparison across datasets, we use a filtered subset of MMLU containing only open-ended questions [myrzakhan2024openllmleaderboardmultichoiceopenstylequestions].
We select these datasets because they represent two distinct task types: knowledge-seeking (TriviaQA) and reasoning-heavy (MMLU). Prior work has found semantic dispersion to be a useful proxy for language model uncertainty in the former task type and less so in the latter [xiong2024efficient]; as such MMLU provides a more challenging setting for evaluating SSD, although in practice it reflects a mixture of both task types.
We evaluate across five model families: Mistral (Ministral-8B-Instruct), Llama-3 (3.1-8B-Instruct, 3.2-3B-Instruct), Qwen (Qwen3-4B-Instruct-2507, Qwen3-8B), Gemma (Gemma-3-4B) and SmolLM (SmolLM3-3B) [grattafiori2024llama, mistralai_ministraux_2024, yang2025qwen3technicalreport, gemmateam2025gemma3technicalreport, bakouch2025smollm3]. These models span multiple architectural families, parameter scales (3B-8B), and training pipelines, allowing us to assess the robustness and generality of SSD across diverse backbone designs rather than tailoring it to a single model.
@@
For all experiments, we use Qwen3-8B as the oracle ‘judge’ model to determine the correctness of generated answers given a reference answer [zheng2023judge]. For each prompt and model, we generate $S=32$ stochastic samples at temperature $T=1$. We additionally sample a single ‘default’ answer at $T=0.1$, which is provided to the judge model to obtain the hallucination label.
@@
For answer embeddings, we use EmbeddingGemma [vera2025embeddinggemmapowerfullightweighttext], with the embedding truncated to a dimension of $d_{z}=128$. For hallucination prediction, we further reduce dimensionality using PCA to $d_{z}^{\prime}=16$, following ablation studies detailed in Appendix C. For likelihood-based OOD verification, we set $d_{z}^{\prime}=32$.
4.2 Prompt-Level Hallucination Prediction
Table 1: Hallucination prediction trade-offs. AUROC is reported per dataset as a macro-average (mean $\pm$ std) across models. Latency summarizes inference-time requirements (single-pass vs. $S$ sampled generations, with entailment-based SE additionally requiring $S^{2}$ natural language inference (NLI) comparisons). Target cost denotes the computation required to construct supervision targets during training; $C_{CL}$ is the cost of acquiring correctness labels. Only SSD produces an explicit predictive density, enabling post-generation likelihood scoring and latent sampling in addition to pre-generation uncertainty.
Method TriviaQA $\uparrow$ MMLU $\uparrow$ Latency $\downarrow$ Target cost $\downarrow$ Predictive density
SE 0.804 $\pm$ 0.035 0.663 $\pm$ 0.040 $S$ samples + $S^{2}$ NLI $\times$
SEP 0.759 $\pm$ 0.026 0.628 $\pm$ 0.029 Single-pass $S$ samples + $S^{2}$ NLI $\times$
PCP 0.771 $\pm$ 0.020 0.679 $\pm$ 0.019 Single-pass $C_{\text{CL}}$ $\times$
TD 0.699 $\pm$ 0.027 0.647 $\pm$ 0.023 $S$ samples $\times$
SSD 0.708 $\pm$ 0.024 0.623 $\pm$ 0.025 Single-pass $S$ samples + $S\,\Phi$ $\checkmark$
Table 2: Hallucination prediction, comparison with teacher dispersion. TD computes dispersion from $S=32$ sampled semantic embeddings at inference time, while SSD predicts a prompt-conditioned semantic density and scores dispersion analytically. We report hallucination prediction AUROC/AUPRC in % as the mean over 1,000 bootstrap resamples of the test set, with standard deviation as subscript; bold indicates the better of TD vs. SSD per metric and dataset.
AUROC AUPRC
Model TD SSD TD SSD
TriviaQA
Qwen3 8B 68.31.6 67.71.8 52.42.6 56.92.8
Qwen3 4B 73.21.6 69.91.6 59.72.4 61.92.5
Llama 3.1 8B 69.61.7 68.71.9 44.42.7 51.92.9
Llama 3.2 3B 66.61.7 73.71.6 55.82.6 65.52.5
Ministral 8B 74.31.7 74.91.6 57.92.8 59.92.8
SmolLM3 3B 69.91.6 71.11.6 57.12.5 62.42.5
Gemma 3 4B 67.61.6 70.01.7 63.12.3 66.12.5
MMLU
Qwen3 8B 63.51.7 57.11.8 61.92.3 55.92.3
Qwen3 4B 63.11.8 64.91.7 58.52.4 59.42.5
Llama 3.1 8B 69.01.6 63.81.7 70.92.1 62.72.4
Llama 3.2 3B 67.11.8 63.81.7 74.02.0 74.51.9
Ministral 8B 65.11.8 63.61.8 71.42.1 71.72.1
SmolLM3 3B 62.31.8 62.01.8 64.62.2 63.02.3
Gemma 3 4B 62.81.7 60.71.8 71.12.0 69.22.1
@@ 4.2 Prompt-Level Hallucination Prediction
We first evaluate whether the distilled semantic density provides a reliable pre-generation hallucination signal. Specifically, we treat the entropy of the predicted distribution as a prompt-level risk score: higher entropy corresponds to greater semantic dispersion and increased hallucination risk. We report AUROC and AUPRC against correctness labels determined using an LLM-as-judge framework.
@@
We compare the student’s dispersion score to the standard deviation of the teacher’s $S=32$ semantic embeddings, which we term teacher dispersion (TD). Table 2 reports results across models and datasets. On the TriviaQA dataset, despite requiring no inference-time sampling, SSD outperforms TD on four of the seven models investigated in terms of AUROC and on all seven in AUPRC. We attribute this in part to a smoothing effect: TD is a finite-sample statistic and assigns zero uncertainty to prompts where the model consistently produces the same answer, even when that answer is incorrect. By contrast, SSD learns a continuous mapping from prompt representations to semantic uncertainty, enabling differentiated risk estimates in such degenerate cases, which likely contributes to its stronger AUPRC.
On the MMLU dataset, SSD is a relatively weaker performer compared to the sampled teacher signal, which we attribute to the task being less suited for the semantic dispersion uncertainty proxy. SSD still surpasses TD in three out of seven models from an AUPRC perspective. Overall, these results demonstrate that SSD successfully distils sampling-based semantic dispersion into a single-pass density estimator. Despite requiring no inference-time sampling, the student recovers much of the teacher’s uncertainty signal, and in some cases improves upon the finite-sample baseline.
@@
We compare SSD against three widely used hallucination prediction methods: (1) Probability of Correctness Probe (PCP), a logistic regression classifier trained on the prompt representation to predict correctness. Originally proposed as ‘P(I know)’ by kadavath2022languagemodelsmostlyknow, this method requires externally provided correctness labels. (2) Semantic Entropy (SE), a sampling-based approach that measures semantic dispersion by clustering $S$ completions using a natural language inference (NLI) model to assess mutual entailment [farquhar_detecting_2024]. (3) Semantic Entropy Probe (SEP), a lightweight probe with the same architecture as PCP, trained to regress a binarized SE target, and inheriting SE’s sampling and NLI-based target construction cost [kossen2024semanticentropyprobesrobust].
@@
To provide a fuller comparison of methods, we report hallucination prediction performance alongside method characteristics: inference latency, training supervision target construction cost, and whether a method supports post-generation signals in addition to pre-generation uncertainty. Table 1 summarizes these trade-offs and reports macro-average AUROC across model families; full per-model AUROC/AUPRC tables are provided in Appendix F.
Overall, SE achieves the highest AUROC on TriviaQA, whilst PCP is the strongest performer on MMLU, consistent with prior work [xiong2024efficient]. While SE and PCP achieve the best performance in their respective regimes, these methods either incur significant sampling cost (SE) or reduce uncertainty to a supervised scalar classification problem (PCP). SSD occupies a distinct regime: it matches probe-level inference latency, avoids entailment-based target construction during training, and preserves an explicit predictive density. Although its pre-generation hallucination prediction performance is slightly weaker than specialized probes, it uniquely provides post-generation likelihood signals derived from the same model.
We further observe that the SSD performance is model-dependent. On TriviaQA, for example, it achieves an AUROC of 74.9 for Ministral 8B but only 67.7 for Qwen3 8B (Table A5). Analysis in Appendix B.1 indicates that this variation is driven primarily by the fidelity of the distillation process: for some models, it is particularly difficult to predict the semantic distribution from the prompt representation. We discuss these findings further in Section 5.
4.3 Answer-Level Response Evaluation
Having established that SSD recovers sampling-based dispersion signals, we next evaluate the additional capabilities enabled by modelling an explicit predictive density. Unlike scalar uncertainty scores, the density allows us to assess the quality of specific candidate answers by measuring their likelihood under the model’s predicted semantic distribution. We examine this utility in two settings: out-of-domain answer detection (binary classification) and multiple-choice question answering (multi-class selection).
4.3.1 Out-of-domain Answer Detection
@@
We first evaluate the student’s ability to detect mismatched prompt-answer pairs. We frame this as an out-of-domain detection task: for each test prompt $x_{i}$, we pair its representation $\mathbf{h}_{i}$ with (i) the model’s default answer $y_{i}$ and (ii) a random answer $y_{j}$ drawn from a different prompt. Each candidate answer is embedded as $\mathbf{z}=\Phi(y)$ and scored using the student’s log-likelihood $\log q_{\phi}(\mathbf{z}\mid\mathbf{h})$. We report AUROC for distinguishing matched from mismatched pairs.
@@
As shown in Table 3, the SSD log-likelihood is a strong context-verification signal, achieving AUROCs above 0.95 for most base models on both TriviaQA and MMLU. Practically, this could enable SSD to act as a content ‘guard’ [inan2023llamaguardllmbasedinputoutput], for example detecting cache poisoning [chenagentpoison2024] or filtering retrieved content in retrieval-augmented generation [lewisrag2020].
We also evaluate the student likelihood as a post-generation hallucination signal, motivated by prior work showing that probability density in semantic space is strongly predictive of correctness [qiu2024semanticdensity]. While informative, our likelihood signal is weaker than the pre-generation entropy, with AUROC scores of approximately 0.6. Unlike sampling-based density methods, which construct the semantic distribution from multiple generations at inference time, SSD predicts the density from the prompt representation alone. The weaker performance reflects the limits of this single-pass approximation to a sampling-derived density.
Table 3: Out-of-domain answer detection. We report AUROC for two binary classification tasks using the student’s log-likelihood score. Mismatch measures the ability to distinguish the model’s answer to the prompt from a random answer from the dataset. Hallucination measures the ability to distinguish correct model-generated answers from incorrect ones. Values report mean AUROC with bootstrap standard deviation over 1,000 resamples.
Model Mismatch Hallucination
TriviaQA
Qwen3 8B 0.984 $\pm$ 0.002 0.571 $\pm$ 0.019
Qwen3 4B 0.984 $\pm$ 0.002 0.557 $\pm$ 0.018
Llama 3.1 8B 0.989 $\pm$ 0.002 0.579 $\pm$ 0.019
Llama 3.2 3B 0.980 $\pm$ 0.003 0.613 $\pm$ 0.017
Ministral 8B 0.945 $\pm$ 0.005 0.616 $\pm$ 0.017
SmolLM3 3B 0.988 $\pm$ 0.002 0.562 $\pm$ 0.018
Gemma 3 4B 0.895 $\pm$ 0.007 0.578 $\pm$ 0.018
MMLU
Qwen3 8B 0.975 $\pm$ 0.003 0.526 $\pm$ 0.018
Qwen3 4B 0.960 $\pm$ 0.004 0.566 $\pm$ 0.019
Llama 3.1 8B 0.981 $\pm$ 0.003 0.576 $\pm$ 0.018
Llama 3.2 3B 0.942 $\pm$ 0.005 0.519 $\pm$ 0.019
Ministral 8B 0.966 $\pm$ 0.004 0.529 $\pm$ 0.019
SmolLM3 3B 0.985 $\pm$ 0.003 0.528 $\pm$ 0.018
Gemma 3 4B 0.923 $\pm$ 0.006 0.548 $\pm$ 0.019
4.3.2 Multiple-choice Answer Selection
We next evaluate whether the student’s predicted density can be used to identify the correct answer in a multiple-choice setting. We utilize the open-ended subset of MMLU provided by myrzakhan2024openllmleaderboardmultichoiceopenstylequestions, which filters for questions that are self-contained and do not require the context of multiple-choice options to answer. The student is trained using only open-ended completions to prompts in the training set. For the test set, we take the four provided options ($A,B,C,D$) for each question, embed them using $\Phi$, and rank them according to their log-likelihood $\log q_{\phi}(\mathbf{z}^{\prime}\mid\mathbf{h})$ under the student mixture.
Table 4 reports the accuracy of selecting the highest-likelihood option (Top-1) alongside the base model’s open-ended answer accuracy as assessed by an LLM-as-judge. SSD consistently exceeds the 0.25 random baseline, and for our best performing models (Llama 3.2 and Ministral) approaches the base model’s open-ended accuracy.
The open-ended accuracy of the base model provides a rough estimate of its dataset-specific knowledge. Despite being trained only on sampled open-ended completions and without multiple-choice supervision, the student approaches this level of performance when provided with the options, suggesting that the distilled semantic density captures meaningful structure in the model’s answer distribution.
Table 4: Multiple-choice answer selection. We evaluate the ability of the SSD student - trained solely on open-ended generations - to identify the correct answer in a multiple-choice setting using an open-ended subset of MMLU. We report Top-1 accuracy, where the option with the highest predicted likelihood is selected (random baseline =0.25=0.25). Open-Ended accuracy denotes the base model’s accuracy when generating answers directly, as determined by an LLM-as-judge.
Model SSD Top-1 Open-Ended
Qwen3 8B 0.327 $\pm$ 0.015 0.515 $\pm$ 0.015
Qwen3 4B 0.312 $\pm$ 0.014 0.545 $\pm$ 0.016
Llama 3.1 8B 0.339 $\pm$ 0.015 0.483 $\pm$ 0.015
Llama 3.2 3B 0.346 $\pm$ 0.015 0.358 $\pm$ 0.015
Ministral 8B 0.363 $\pm$ 0.015 0.396 $\pm$ 0.015
SmolLM3 3B 0.332 $\pm$ 0.015 0.466 $\pm$ 0.016
Gemma 3 4B 0.292 $\pm$ 0.014 0.399 $\pm$ 0.015
Summary of experimental findings.
Taken together, the prompt-level and answer-level evaluations demonstrate that SSD’s value lies not in a particular scalar uncertainty score, but in exposing an explicit semantic belief state from which multiple reliability signals can be derived. By modelling a prompt-conditioned distribution rather than a scalar proxy, SSD unifies pre-generation risk estimation and post-generation verification within a single, low-latency framework. The performance variation observed across base LLMs points to differences in how semantic uncertainty is encoded in internal activations.
5 Discussion & Future Work
The experimental results highlight both the strengths and current limitations of SSD. While modelling a full semantic density enables richer post-generation capabilities than scalar probes, its effectiveness depends on how faithfully the student can reconstruct the teacher’s sampled distribution from prompt representations alone. We now discuss these limitations and outline directions for strengthening the distributional framework.